OpenStudy (conqueror):

@shrutipande9 @thomaster @AravindG @myininaya @ganeshie8

OpenStudy (conqueror):

Me and my friend decided to stop people from posting test questions and people giving direct answers, and my friend just got a suspension for doing that.

OpenStudy (aravindg):

What was your friend's name?

OpenStudy (camerondoherty):

You're supposed to go to abuse.openstudy.com

OpenStudy (conqueror):

LearningisAwesome

OpenStudy (camerondoherty):

i think

OpenStudy (aravindg):

Let me know the time of suspension and possible reasons.

OpenStudy (conqueror):

24 hours, can't find any possible solutions. Didn't you say you were glad that we were doing that stuff? @AravindG .-.

OpenStudy (shrutipande9):

@Conqueror i already replied in the math section.

OpenStudy (aravindg):

^I think you already have an answer.

OpenStudy (conqueror):

Oh didn't see yet @shrutipande9 Hold on

OpenStudy (conqueror):

@ArmedWithMadden What was the reason they said you got suspended?

thomaster (thomaster):

24 hours is usually an automatic suspension after too many people report you.

OpenStudy (anonymous):

"Too many abuse reports" - to be clear I haven't hit the report abuse button once all day.

thomaster (thomaster):

yes too many abuse reports means that you were suspended automatically by the server. Happens when an x amount of users report your messages/ answers/ chats.

OpenStudy (conqueror):

So people reported @ArmedWithMadden ?

OpenStudy (conqueror):

Only because he was telling them to obey the Code of Conduct??

OpenStudy (anonymous):

Well they hit that because they were mad at me for giving links to the CoC, not because I did something wrong.

OpenStudy (anonymous):

Can a mod lift a suspension? T-T

OpenStudy (camerondoherty):

Why don't you leave the code of conduct and people giving straight answers thing to the mods?

OpenStudy (conqueror):

The mods don't take care of it sometimes, and @AravindG said he was glad we were doing it.

OpenStudy (anonymous):

We're not having this discusion @camerondoherty

OpenStudy (anonymous):

Had it wayyyyy too many times.

OpenStudy (anonymous):

But really I just told people the CoC very politely and now I'm suspended, so can a mod please lift it or something? T-T

OpenStudy (aravindg):

I will have to check with the moderator log. Please show patience. Thanks!

OpenStudy (anonymous):

What exactly does that mean? :/ not impatient just really confused on what that means.

OpenStudy (anonymous):

@AravindG

OpenStudy (armyrangers):

@ArmedWithMadden aren't you braking Code of conduct right now

myininaya (myininaya):

The mods take care of it when we see it. So when people report it, it helps bring it to our attention @Conqueror A moderator log is a log of actions that have been recorded. He is looking through it. He will respond in a sec.

OpenStudy (aravindg):

It meant I needed some time to check it.

OpenStudy (anonymous):

Ohhh okay.

OpenStudy (anonymous):

@camerondoherty that's me. I made another account just to ask why I got suspended, after no reply from the abuse email.

OpenStudy (camerondoherty):

ik thts u

OpenStudy (aravindg):

Okay done! Welcome back @LearningIsAwesome :) Sorry that you got automatically suspended under false reports.

OpenStudy (anonymous):

YAAYYYYYYYYYYYYYYYYYYYYYYYYYYYYYYYYYYYYYYYYY!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

I AM BATSTUDY...

OpenStudy (camerondoherty):

Congrats on regaining your account c:

OpenStudy (anonymous):

@Conqueror and @AravindG are so my favorite people on planet earth. I love you guys :3

OpenStudy (conqueror):

Iron Man does his duty.

OpenStudy (anonymous):

AND SO DOES BATSTUDY

OpenStudy (anonymous):

Wait, shouldn't you be Iron Study?

OpenStudy (abhisar):

Sweet ! \(\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \Large\cdot \quad \cdot\\ \cdot\\ \huge \smile \end{matrix} \right)}}\normalsize \\ \;/\quad \;\;\quad \backslash\)

OpenStudy (camerondoherty):

maybe you should close this now... since your problem was solved...